OpenStudy (anonymous):

what is the equation for the line that passes through points (-1,-4) and (1,4)? write the equation in slope-intercept form.

OpenStudy (anonymous):

@StudyGurl14 @3spooky5me @Catlover5925 @Here_to_Help15 @perl

OpenStudy (anonymous):

@invinceble @Monster1234 @Owlfred @Zarkon

OpenStudy (studygurl14):

Use the slope formula: Where points \((x_1,y_1)\) and \((x_2,y_2)\) are on the same line... \(\large slope=\Large\frac{y_2-y_2}{x_2-x_1}\)

OpenStudy (anonymous):

slope=4 @StudyGurl14

OpenStudy (studygurl14):

Correct. good job. :)

OpenStudy (studygurl14):

Slope-intercept form: y= mx+b where m =slope So... y= 4x + b You need to find the value of b. Do that by plugging in one of the coords into the equation and solving. Remember that (x,y)

OpenStudy (anonymous):

so y=4x+b y=4 times 1 +4 is the equation @StudyGurl14

OpenStudy (studygurl14):

no..not quite Using point (1,4) \(\large y=4x+ b\rightarrow (4) = 4(1) + b \rightarrow 4 = 4 + b\rightarrow b=0\) So... the equation is \(\large y=4x\)

OpenStudy (anonymous):

ok i get it now bc i looked it up so thanks

OpenStudy (studygurl14):

You're welcome. :)